# Chenjie Zhong    Zhipeng Li†,*    Shangzhi Xu Xiaohu Li

### v2

@@ Abstract
Automatically determining the number of intrinsic mode functions (IMFs) and their center frequencies in Variational Mode Decomposition (VMD) still remains an open mathematical challenge, which often relies on heuristic prior settings, trial-and-error strategies or complex-field based methods that lack of theoretical guarantees on the convergence. In this article, we proposed a novel variational method based on real field, focusing on evaluating the supporting baseline of a signal’s spectrum, to further retrieve the intrinsic mode functions. Our method treats automatic extraction of modes as a constrained optimization problem that adversarially maximizes the baseline integral while penalizing its curvature, and transform the problem into iteratively solving a fourth-order boundary value problem via Lagrangian duality. Furthermore, we establish the rigorous mathematical proof on the global convergence to our algorithm based on dual ascent in function space. Comprehensive numerical experiments on artificial and real-world signals including electrocardiogram (ECG) data show that our method can provide accurate estimates of IMFs and center frequencies, and comparison with methods like Successive VMD also shows our advantages in avoiding redundant modes while ensuring the recovery of necessary components, indicating that we have provided a robust, theoretically grounded initialization routine for VMD.
@@
narrow-banded signals;variational optimization;Intrinsic Mode Function
{MSCcodes}
65K10, 65L10, 68U01, 46N10, 34B05
@@ 1 Introduction
(6) $\hat{u}_{k}^{n+1}=\underset{u_{k},\,\hat{u}_{k}=\hat{u}_{k}^{T}}{\arg\min}\left\{\int_{0}^{\infty}4\alpha(\omega-\omega_{k})^{2}|\hat{u}_{k}(\omega)|^{2}d\omega+2\int\left(\hat{f}-\sum\hat{u}_{i}+\frac{\hat{\lambda}}{2}\right)^{2}d\omega\right\}$
@@
(8) $\omega_{k}^{n+1}=\frac{\displaystyle\int_{0}^{\infty}\omega\left|\hat{u}_{k}(\omega)\right|^{2}d\omega}{\displaystyle\int_{0}^{\infty}\left|\hat{u}_{k}(\omega)\right|^{2}d\omega}$
Since its establishment, VMD is always playing an important role in modern signal decomposition in many fields[14, 21, 9, 7, 16, 15, 1]. Although VMD has provided flexibility in decomposing signals, there are factors that hinder its further applications in that the number of intrinsic mode function must be determined manually as a prior in coordination with penalty factor without a common principle[19, 10, 18, 17]. In order to alleviate the problem, the subsequent research either needs the knowledge to the range of the number of modes ,[11] retrieves the IMFs recursively ,[12] or highly dependent on a predefined window-size to retrieve the peak and valley of spectra [6]. All these methods either require prior or rely on schemes devoid of convergence guarantees, making the results contains spurious or missed modes. The absence of a well-posed, convergent paradigm for automatically finding the number of IMFs is a crucial factor that hinders the development of VMD.
To close this gap, we introduce a fundamentally different approach for finding the IMFs. Instead of retrieving modes in an trial-and-error way or recursively extracting modes, we focus on the spectrum amplitude and reformulate the problem by introducing supporting baseline, a lower-enveloping curve of a signal’s spectrum which can support the spectrum as tightly as possible while keeping itself as smooth as possible, and transform the problem for finding IMFs into first evaluating the supporting baseline by adversarially maximizing its integral while penalizing its curvature, then retrieving the separated peaks of the original spectrum above the supporting baseline.
Our main contribution can be summarized as 4 folds. First, we give a new perspective for determining the number of IMFs by introducing supporting baseline of spectrum and solve the problem in real field. Second, we derive its optimality conditions, establishing a rigorous equivalence between this variational problem and a fourth-order boundary value problem, and establish rigorous mathematical proof for the global convergence to the course of dual ascent iteration in function space. Thirdly, we develop an efficient iteration scheme for the algorithm by converting the corresponding difference equation and boundary conditions into matrix format and expanding compatible rule to the component-wise multiplication between matrices and vectors for Hadamard product. Finally, we give experiments on both artificial and real signals, and comparison with SVMD to show the robustness of our algorithm and experiments show that our method can quickly get accurate estimation about the number of IMFs and the corresponding initial center frequencies as prior to the following VMD procedure.
@@ 2 Methodology and Convergence Proof
In this section we introduce our method for estimating the number of IMF modes as well as center frequencies of each band. We first introduce the background of the problem, and then we propose the concept of Supporting Baseline of a Function, finally we describe our method for finding the baseline and give the convergence proof for our method.
@@
Figure 1: Illustration to the concept of supporting baseline: In (a)-(c) the solid line is $y=10\cos\pi x-10\cos 5\pi x+x^{2}$. The dashed line in (a) is a proper supporting baseline since it actually fits the lower bound of the solid line as well as ignoring high frequent details of the solid line. The dashed line in (b) has a gap to the bottom of the solid line so it is not suitable to be a good supporting baseline. In contrast, the dashed line in (c) fits too tightly to the lower bound of the solid line, capturing too much detail of the solid line so that it can be reckoned as the lower envelope rather than a proper supporting baseline.
2.1 Supporting Baseline of a Function
@@
Without loss of generality we first introduce the Supporting Baseline of a function. For any function $f(x)>0$, the supporting baseline is a curve that tightly fits the lower bound of the function while reflects the overall trend of the lower bound. Fig. 1 gives an intuitive illustration about the conception of supporting baseline. As illustrated in Fig. 1, intuitively the supporting baseline should be tightly close to the lower bound of the function, approaching the infimum of functions as close as possible as well as maintain its own smoothness in order to capture the global trend of the lower bound rather than too much details, as capturing too much details will in turn reduce the robustness for separating the modes since there might be some noise added to the signal in either time or frequency domain.
2.2 Variational Methods for Finding the Supporting Baseline
@@
To solve the problem for finding the supporting baseline, we use optimization method mainly based on variation and convex optimization[8, 4]. Let $f(x)\geq 0$ be a known function, the problem of finding the supporting baseline curve of $f(x)$ is equivalent to finding a function $g(x)$ that tightly approaching the point-wise infimum of $f(x)$ while maintaining the global trend of its lower bound. This can be viewed as the following optimization problem:
@@
where 0 stands for the left end and $n$ the right end of the curve. To solve the boundary value problem of ODE, we use Finite Difference Method.[20] Note that for a differential function $g(x)$, it’s derivative of first to fourth order can be expressed as the linear combination of $g(x)$ itself in the corresponding differential format and we can take similar operation to the function $\alpha(x)$ (See Supplementary Materials A), so that different orders of derivative with respect to $\alpha(x)$ and $g(x)$ can be expressed as the linear combination of themselves, one can easily transform (19) into the matrix-form as follows:
@@
Algorithm 1 The Key Steps for Evaluating Support Baseline
@@ 2.3.3 Convergence in Dual Space
These assumptions are standard for analyzing variational problems involving curvature penalties. In practice, for discrete and potentially noisy spectral data, $f(x)$ can be considered as a smooth interpolation of the observed points. Crucially, the core object of interest—the supporting baseline—is intrinsically a smooth function due to the curvature penalty term in the objective. Therefore, the analysis under these smoothness assumptions captures the essential behavior of the algorithm.
@@
CASE II: If the optimum of $\mathcal{G}$ forms a convex set, and during the iteration the sequence $(\lambda_{k}(x),\mu_{k}(x))$ is approaching and never fall into the optimal set, then from the middle parts of (51) and (45) we know that the value of $\mathcal{G}$ is approaching the minima and the distance between adjacent term of the sequence $(\lambda_{k}(x),\mu_{k}(x))$ and $(\lambda_{k+1}(x),\mu_{k+1}(x))$ is yielding zero. Thus we know in this case the iterating sequence in dual space can always terminate to the point as close as possible to one of the optimum, although the classical convergent behaviour for the sequence in this case might not be guaranteed.
@@
CASE IV: If the optimum of $\mathcal{G}$ forms a convex set, and the iterating sequence has approached to one of the optimum $(\lambda^{*}_{1}(x),\mu^{*}_{1}(x))$ early during some iterating step, then since in the convex set there are at least another optima $(\lambda^{*}_{2}(x),\mu^{*}_{2}(x))$, then we have
(53) $\mathcal{G}(m\lambda^{*}_{1}(x)$ $+(1-m)\lambda^{*}_{2}(x),m\mu^{*}_{1}(x)+(1-m)\mu^{*}_{2}(x))$
$=\mathcal{G}(\lambda^{*}_{1}(x),\mu^{*}_{1}(x))=\mathcal{G}(\lambda^{*}_{2}(x),\mu^{*}_{2}(x)),m\in[0,1]$
since the direction of the line connecting the median point and the initial point in the first-order Taylor expansion formula is consistent with $\left[\begin{array}[]{c}\lambda^{*}_{2}(x)-\lambda^{*}_{1}(x)\\ \mu^{*}_{2}(x)-\mu^{*}_{1}(x)\end{array}\right]$, so we have
(54) $D\mathcal{G}_{(\lambda_{k}(x),\mu_{k}(x))}(\lambda(x),\mu(x))\left[\begin{array}[]{c}\lambda^{*}_{2}(x)-\lambda^{*}_{1}(x)\\ \mu^{*}_{2}(x)-\mu^{*}_{1}(x)\end{array}\right]=0$
thus the variation direction is always orthogonal to any lines connected between different optima, as a consequence, when projected again during the next iteration, $(\lambda_{k+1}(x),\mu_{k+1}(x))$ must overlap with $(\lambda_{k}(x),\mu_{k}(x))$, so we have $(\lambda_{k+1}(x),\mu_{k+1}(x))=(\lambda_{k}(x),\mu_{k}(x))$.
@@ 2.3.4 Convergence in Primal Space
As the solution of (10) is continuously dependent on its coefficients, so we have $\lim_{k\rightarrow+\infty}g_{k}(x)=g^{*}(x)$ when $\{(\lambda_{k}(x),\mu_{k}(x))\}$ when in Case I, III and IV, and $g_{k}(x)$ is approaching $g^{*}(x)$ when $g_{k+1}(x)-g_{k}(x)||$ is approaching zero in Case II respectively. In all cases we can terminate algorithm and guarantee the $\{g_{k}(x)\}$ terminates to one of the optimum as close as possible.
@@ 3 Implementation
In this chapter, we focus on the detail about the implementation and numerical scheme about the algorithm for evaluating problem (10). Our workflow consists of 4 interconnected stages: (1) We discretize (19) by using Finite Difference Method,[20] transforming it into a linear system. (2) Crucially, by expressing the derivations of both solution $g(x)$ and the weight function $\alpha(x)$ as linear combination of their nodal values, we recast the entire ODE into a compact matrix form by introducing component-wise multiplication with a mild extension of Hadamard production. By imposing well-posed boundary conditions, we make the coefficient matrices full-rank and the whole problem invertible so that the evaluation can be significantly accelerated by modern software toolkits via matrix operation. (3) To further accelerate the convergence and make algorithm effective when handling high or low-passed signals, a smooth extrapolation technique is applied to the input spectrum at the both ends. (4) Finally, after obtaining the optimal baseline, a statistically grounded threshold determined via kernel density estimation (KDE), is employed to automatically identify significant spectral peaks, thereby completing the mode extraction. The following subsections detail each of these components.

### v3

@@ Abstract
Automatically determining the number of intrinsic mode functions (IMFs) and their center frequencies in Variational Mode Decomposition (VMD) remains an open mathematical challenge. Existing methods often rely on heuristic prior settings, trial-and-error strategies, or recursive extraction procedures that are computationally inefficient and prone to error accumulation, and generally lack theoretical guarantees on convergence. In this article, we propose a novel variational framework that endogenously determines the number of modes. We observe that any curve lying below the spectral amplitude naturally divides the spectral domain into connected regions where the spectrum rises above the curve, the number of which defines the modal number K[g]K[g]—a topological functional induced by the cutting curve. Since $K[g]$ is discontinuous and intractable for direct optimization, we instead seek the optimal cutting curve itself as a continuous variational surrogate: the optimal curve best separates distinct spectral peaks into individual connected regions above it, while merging noise-induced fragments below. This surrogate adversarially maximizes the integral of $g$ (encouraging it to rise and support significant peaks) while penalizing its curvature (suppressing excessive undulations that would fragment the spectrum), transforming the problem into iteratively solving a fourth-order boundary value problem via Lagrangian duality. We establish a rigorous proof of global convergence for the dual ascent algorithm in function space. Comprehensive numerical experiments on artificial and real-world signals including electrocardiogram (ECG) data show that our method can provide accurate estimates of IMFs and center frequencies, and comparison with methods like Successive VMD also shows our advantages in avoiding redundant modes while ensuring the recovery of necessary components, indicating that we have provided a robust, theoretically grounded initialization routine for VMD.
@@
Variational Optimization;Convex Optimization;Intrinsic Mode Function;Cutting Curve;Variational Mode Decomposition;Signal Processing
@@ 1 Introduction
(6) $\hat{u}_{k}^{n+1}=\underset{u_{k},\,\hat{u}_{k}=\hat{u}_{k}^{T}}{\arg\min}\left\{\int_{0}^{\infty}4\alpha(\omega-\omega_{k})^{2}|\hat{u}_{k}(\omega)|^{2}+2\left(\hat{f}-\sum\hat{u}_{i}+\frac{\hat{\lambda}}{2}\right)^{2}\text{d}\omega\right\}$
@@
(8) $\omega_{k}^{n+1}=\frac{\displaystyle\int_{0}^{\infty}\omega\left|\hat{u}_{k}(\omega)\right|^{2}\text{d}\omega}{\displaystyle\int_{0}^{\infty}\left|\hat{u}_{k}(\omega)\right|^{2}\text{d}\omega}$
Since its establishment, VMD is always playing an important role in modern signal decomposition in many fields[14, 21, 9, 7, 16, 15, 1]. Although VMD has provided flexibility in decomposing signals, there are factors that hinder its further applications in that the number of intrinsic mode function must be determined manually as a prior in coordination with penalty factor without a common principle[19, 10, 18, 17]. In order to alleviate the problem, the subsequent research either needs the knowledge to the range of the number of modes[11], retrieves the IMFs recursively[12], or highly dependent on a predefined window-size to retrieve the peak and valley of spectra [6]. All these methods either require prior or rely on schemes devoid of convergence guarantees, making the results contains spurious or missed modes. The absence of a well-posed, convergent paradigm for automatically finding the number of IMFs is a crucial factor that hinders the development of VMD.
To close this gap, we introduce a fundamentally different approach for finding the IMFs. Instead of retrieving modes in an trial-and-error way or recursively extracting modes, we focus on the spectrum amplitude and reformulate the problem by introducing Cutting Curve, a lower-enveloping curve of a signal’s spectrum which can support the spectrum as tightly as possible while keeping itself as smooth as possible, and transform the problem for finding IMFs into first evaluating the Cutting Curve by adversarially maximizing its integral while penalizing its curvature, then retrieving the separated peaks of the original spectrum above the Cutting Curve.
Our main contribution can be summarized as 4 folds. First, we give a new perspective for determining the number of IMFs by introducing Cutting Curve of spectrum and solve the problem in real field. Second, we derive its optimality conditions, establishing a rigorous equivalence between this variational problem and a fourth-order boundary value problem, and establish rigorous mathematical proof for the global convergence to the course of dual ascent iteration in function space. Thirdly, we develop an efficient iteration scheme for the algorithm by converting the corresponding difference equation and boundary conditions into matrix format and expanding compatible rule to the component-wise multiplication between matrices and vectors for Hadamard product. Finally, we give experiments on both artificial and real signals, and comparison with SVMD to show the robustness of our algorithm and experiments show that our method can quickly get accurate estimation about the number of IMFs and the corresponding initial center frequencies as prior to the following VMD procedure.
@@ 2 Methodology and Convergence Proof
According to the theory of Variational Mode Decomposition, each intrinsic mode function should manifest in the spectral domain as a prominent convex bulge, or alternatively as a cluster of adjacent minor bulges. In an ideal, flat spectrum, identifying such bulges would be straightforward. However, real-world spectra invariably exhibit an underlying trend—a slowly varying floor upon which the modal peaks sit. This trend obscures the boundaries between modes, rendering the extraction of bulges far from trivial.
In this section, we develop a principled framework to address this difficulty. Our strategy proceeds as follows. First, we formalize the intuitive conditions that an optimal modal partition should satisfy: each detected mode must accumulate sufficient residual energy with respect to the trend of the spectrum; adjacent weak peaks that individually lack saliency should be merged; and distinct modes should be separated by sufficiently deep valleys. Second, let $f(x)$ be the spectrum of any investigated signal, we recognize that any curve $g(x)$ lying below $f(x)$ will introduce a partition with respect to the frequency-axis into connected intervals where f(x)g(x)>0f(x)-g(x)>0. The number of each connected intervals defines the modal number $K(g(x))$ which is a general function of $g(x)$, and the problem of determining the optimal number $K$ is then recast to selecting an optimal curve $g(x)$ to cut the spectrum, or the optimal Cutting Curve $g^{*}(x)$.
Since $K(g(x))$ is a discontinuous functional that is hard to be optimized, we construct an surrogate continuous functional which can automatically yields a cutting curve satisfying our assumption that mentioned above. The surrogate functional defines a conditional variational problem that adversarially maximizing the integral of the curve and minimizing the curvature. By finding the optimal curve defined by the surrogate function satisfying the assumption of optimal $K$, one can evaluate the number of modes in a continuous functional framework, which is much more easier than imposed in the discontinuous circumstances.
Additionally, we can prove that the optimization of the surrogate functional is a convex problem which can be transformed into an variational problem, and be transformed into a fourth-ordered boundary-value-problem, and a projected dual-ascent algorithm is developed to get the optimal solution. We can also establish the global convergence to the unique minimizer in the function space. The remainder of the part provides the mathematical detail of the proposed method.
2.1 Modal Number as a Functional of the Cutting Curve
In this section we introduce Cutting Curve of a Spectrum.
Definition 1: Cutting Curve and Support Interval Sets
Let $f(x)\geq 0$ be any amplitude spectrum of a signal defined on a closed interval $\Omega\subset\mathbb{R}$, for any continuous function $g(x)\in C(\Omega)$ satisfying $0\leq g(x)\leq f(x)$, we define the supporting intervals $\mathcal{S}(g(x))$ as
(9) $\mathcal{S}(g(x))=\{x\in\Omega|0\leq g(x)\leq f(x)\}$
and $g(x)$ as a cutting curve, then $\mathcal{S}(g(x))$ can be uniquely decomposed into a series countably many disjoint open sets
(10) $\mathcal{S}(g(x))=\bigcup_{k=1}^{K(g(x))}C_{k}$
where
(11) $C_{i}\cap C_{j}=\varnothing,1\leq i\leq j\leq K(g(x))$
Then for any cutting function $g(x)$, $K$ is the uniquely functional with respect to $g(x)$ which can be defined as the number of connected components of the set $\mathcal{S}$.
Definition 2: Optimal Modal Number $K^{*}$
The best number of modal should satisfy the following conditions:
(a) Each component $C_{k}$ of $\mathcal{S}(g(x))$ should support sufficient energy of the spectrum, that is the residual $\int_{C_{k}}f(x)-g(x)\text{d}x$ should be sufficiently significant, with respect to the whole trend of the spectrum.
(b) Between distinct modal components, the cutting curve should remain sufficiently smooth, avoiding spurious oscillations induced by noise or minor fluctuations.
(c) Adjacent weak peaks lacking individual saliency should not constitute separate modes. They are to be either aggregated into a combined mode or, where a dominant peak is present, assimilated as subsidiary ripples of that principal mode.
As for assumption (a), one should find a curve that approaches the infinitum $f(x)$, yet maximizing the residual $\int f(x)-g(x)\text{d}x$, which in turn encourage the evaluated curve to isolate different modes. However, purely encouraging the curve to approach $f(x)$ everywhere will result in nothing to be retrieved, so that the curve should remain at the bottom of $f(x)$ at the interval where peaks exist. By considering (b) and (c), the penalty to the curve is a good choice to satisfy the assumption. Thus by encouraging the integral of $g(x)$ and discouraging the curvature of $g(x)$, evaluating the optimal $g^{*}(x)$ is implicitly evaluating the best mode number $K$ under our basic assumption.
However, directly optimization to $K(g(x))$ is difficult since this is not a continuous functional with respect to $g(x)$, so in the following sections we instead optimize $g(x)$ to approach the target.
@@ 2.1 Modal Number as a Functional of the Cutting Curve
Figure 1: Illustration to the concept of best Cutting Curve: In (a)-(c) the solid line is $y=10\cos\pi x-10\cos 5\pi x+x^{2}$. The dashed line in (a) is a proper Cutting Curve since it actually fits the lower bound of the solid line as well as ignoring high frequent details of the solid line. The dashed line in (b) has a gap to the bottom of the solid line so it is not suitable to be a good cutting curve. In contrast, the dashed line in (c) fits too tightly to the lower bound of the solid line, capturing too much detail of the solid line so that it can be reckoned as the lower envelope rather than a proper Cutting Curve.
2.2 Best Cutting Curve of a Function
@@
Without loss of generality we first introduce the Cutting Curve of a function. For any function $f(x)>0$, the best Cutting Curve is a curve that tightly fits the lower bound of the function while reflects the overall trend of the lower bound. Fig.1 gives an intuitive illustration about the conception of Cutting Curve. As illustrated in Fig.1, intuitively the Cutting Curve should be tightly close to the lower bound of the function, approaching the infimum of functions as close as possible as well as maintain its own smoothness in order to capture the global trend of the lower bound rather than too much details, as capturing too much details will in turn reduce the robustness for separating the modes since there might be some noise added to the signal in either time or frequency domain.
2.3 Variational Methods for Finding the Best Cutting Curve
@@
To solve the problem for finding the best Cutting Curve, we use optimization method mainly based on variation and convex optimization[8, 4]. Let $f(x)\geq 0$ be a known function, the problem of finding the Cutting Curve curve of $f(x)$ is equivalent to finding a function $g(x)$ that tightly approaching the point-wise infimum of $f(x)$ while maintaining the global trend of its lower bound. This can be viewed as the following optimization problem:
@@
where 0 stands for the left end and $n$ the right end of the curve. To solve the boundary value problem of ODE, we use Finite Difference Method[20]. Note that for a differential function $g(x)$, it’s derivative of first to fourth order can be expressed as the linear combination of $g(x)$ itself in the corresponding differential format and we can take similar operation to the function $\alpha(x)$ (See Supplementary Materials A), so that different orders of derivative with respect to $\alpha(x)$ and $g(x)$ can be expressed as the linear combination of themselves, one can easily transform (22) into the matrix-form as follows:
@@
Algorithm 1 The Key Steps for Evaluating Optimal Cutting Curve
@@ 2.4.3 Convergence in Dual Space
These assumptions are standard for analyzing variational problems involving curvature penalties. In practice, for discrete and potentially noisy spectral data, $f(x)$ can be considered as a smooth interpolation of the observed points. Crucially, the core object of interest—the Cutting Curve—is intrinsically a smooth function due to the curvature penalty term in the objective. Therefore, the analysis under these smoothness assumptions captures the essential behavior of the algorithm.
@@
(40) $\begin{bmatrix}\tilde{\lambda}_{k+1}(x)\\ \tilde{\mu}_{k+1}(x)\end{bmatrix}=\begin{bmatrix}\tilde{\lambda}_{k}(x)\\ \tilde{\mu}_{k}(x)\end{bmatrix}+\eta D\mathcal{G}(\lambda_{k}(x),\mu_{k}(x))$
as iteration step and rearrange, we have
@@
CASE II: If the optimum of $\mathcal{G}$ forms a convex set, and during the iteration the sequence $(\lambda_{k}(x),\mu_{k}(x))$ is approaching and never fall into the optimal set, then from the middle parts of (55) and (49) we know that the value of $\mathcal{G}$ is approaching the maxima and the distance between adjacent term of the sequence $(\lambda_{k}(x),\mu_{k}(x))$ and $(\lambda_{k+1}(x),\mu_{k+1}(x))$ is yielding zero. Thus we know in this case the iterating sequence in dual space can always terminate to the point as close as possible to one of the optimum, and the iterating sequence is of asymptotic regularity.
@@
CASE IV: If the optimum of $\mathcal{G}$ forms a convex set, and the iterating sequence has approached to one of the optimum $(\lambda^{*}_{1}(x),\mu^{*}_{1}(x))$ early during some iterating step, then in this case the left side of (49) must be zero since under reasonable step size the value of functional $\mathcal{G}$ cannot decrease, so that the middle part of (49) is zero, indicating the iterating sequences will stay still.
@@ 2.4.4 Convergence in Primal Space
As the solution of (13) is continuously dependent on its coefficients, so we have $\lim_{k\rightarrow+\infty}g_{k}(x)=g^{*}(x)$ when $\{(\lambda_{k}(x),\mu_{k}(x))\}$ when in Case I, III and IV, and $g_{k+1}(x)-g_{k}(x)||$ is approaching zero in Case II respectively. In all cases we can terminate algorithm and guarantee the $\{g_{k}(x)\}$ terminates to one of the optimum as close as possible.
@@ 3 Implementation
In this chapter, we focus on the detail about the implementation and numerical scheme about the algorithm for evaluating problem (13). Our workflow consists of 4 interconnected stages: (1) We discretize (22) by using Finite Difference Method[20], transforming it into a linear system. (2) Crucially, by expressing the derivations of both solution $g(x)$ and the weight function $\alpha(x)$ as linear combination of their nodal values, we recast the entire ODE into a compact matrix form by introducing component-wise multiplication with a mild extension of Hadamard production. By imposing well-posed boundary conditions, we make the coefficient matrices full-rank and the whole problem invertible so that the evaluation can be significantly accelerated by modern software toolkits via matrix operation. (3) To further accelerate the convergence and make algorithm effective when handling high or low-passed signals, a smooth extrapolation technique is applied to the input spectrum at the both ends. (4) Finally, after obtaining the optimal cutting curve, a statistically grounded threshold is determined via kernel density estimation (KDE) on the residual spectrum to filter out background noise. The resulting significant peaks are then aggregated using DBSCAN to merge spatially adjacent minor peaks into a single coherent intrinsic mode, thereby completing the mode extraction. The following subsections detail each of these components.